a train starts from a station with acceleration 0.2 m/s2 on a straight track and then comes to rest after attaining maximum speed on another station due to retardation 0.4m/s2. if total time spent is half an hour, then distance between two stations is {neglect length of train}
1. 216 km
2. 512 km
3. 728 km
4. 1296 km

Grade:

2 Answers

Arun Kumar IIT Delhi
9 years ago
s1=0.5 *0.2 t1^2
s2=vm^2/2*0.4

t1=2*t2

1200,600s

applying kinematic equations
we get s1=144km
s2=72km

a
Arun Kumar IIT Delhi
9 years ago
a